A coil of length 4cm and width 2cm and 45 turns is placed in a uniform field of flux density 0.8T. if the axis of the coil is at right angle to d magnetic field and has a current of 12A, calculate the turning moment on the coil


1
Expert's answer
2021-03-10T17:16:05-0500

Answer

Area of coil


A="Length \times width"

="0.04\times0.02=0.0008m^2"

No. Of turns

"N=45"

Mangnetic field density

B=0.85T

Right angle means

"\theta=90"

Therefore

the turning moment on the coil

"\tau=NIABsin\theta"

"=45\times12\times0.0008\times0.85\times sin90°"

=390.15 N-m